# A non-linear quantum neural network framework for entanglement engineering

### v1

@@ Abstract
Multipartite entanglement is a key resource for quantum technologies, yet its scalable generation in noisy quantum devices remains challenging. Here we propose a low-depth quantum neural network architecture with linear scaling, inspired by memory-enabled photonic components, for variational entanglement engineering. The network incorporates physically motivated non-linear activation functions, enhancing expressivity beyond linear variational circuits at fixed depth. By Monte Carlo sampling over circuit topologies, we identify architectures that efficiently generate highly entangled pure states, approaching the GHZ limit, and demonstrate a clear advantage of non-linear networks up to 20 qubits. For the noisy scenario, we employ the experimentally accessible Meyer–Wallach global entanglement as a surrogate optimization cost and certify entanglement using bipartite negativity. For mixed states of up to ten qubits, the optimized circuits generate substantial entanglement across both symmetric and asymmetric bipartitions. These results establish an experimentally motivated and scalable variational framework for engineering multipartite entanglement on near-term quantum devices, highlighting the combined role of non-linearity and circuit topology.
@@ I Introduction
Entanglement engineering can be formally cast as a graph-optimization problem, in which both interaction structure and parametrization determine the entanglement properties of the resulting quantum state [QuantumExperimentsAndGraphs, EntanglementGenerationKobra]. This perspective has motivated the use of machine-learning techniques, particularly variational optimization and reinforcement learning, in quantum information science [Review_AIforQauntum-Krenn, ReviewAIforQuantumComp-2025, RL_subgraphOptimization]. Such approaches have been successfully applied to the automated design of quantum experiments [VAE_EntanglementExperiments-2022], circuit synthesis [RL4qEntangledQuibits, DiffusionCircuitsynthesis-2024], circuit optimization [RL_circuitOPtimization], and the preparation and engineering of entangled states [RLStatePReparation, RLEntanglementStatePreparation, RL_entanglementEgineering].
@@
Quantum neural networks (QNNs) provide a flexible framework for these tasks by optimizing parametrized quantum circuits using classical learning algorithms [schuld2021supervisedquantummachinelearning]. However, most existing QNN architectures rely on strictly linear parametrizations, limiting their expressive power at fixed circuit depth. Increasing depth can partially alleviate this limitation but leads to higher hardware noise and reduced scalability, making low-depth, linearly scaling architectures particularly attractive for near-term devices [Luo_2023].
Here, we combine statistical learning with photonic quantum hardware by embedding beam-splitter elements, which have a strong experimental track record in entanglement generation [BeamPLsitter_Entanglement_Kumar_2023], into a variational QNN architecture with linear $O(n)$ gate complexity and classical optimization [Goodfellow-et-al-2016]. Compared to more resource-intensive approaches, such as quantum restricted Boltzmann machines [QRBM4Entanglement], the resulting model is simpler to train and scalable without post-selection.
A key ingredient of our approach is the introduction of physically motivated non-linear behavior. While non-linear activation functions are essential for the success of classical neural networks, only a few quantum architectures incorporating non-linearity have been explored so far [QCNN-nonlinear, Reservoir-Spagnolo2022, CVNonlinearQNN]. We draw inspiration from photonic quantum memristors, which exhibit memory-dependent dynamics and support entanglement generation [AlbertoMemristor]. By enforcing a physically inspired non-linear response on the gate parameters, we directly assess the role of activation functions in entanglement engineering [Activactoins-Kunc2024].
The resulting non-linear response naturally leads to sinusoidal activation functions, reflecting the wave nature of the underlying physical processes. This establishes a connection with recent advances in classical machine learning, notably implicit neural representations with periodic activation functions [mildenhall2022neuralfields, SIREN, SIREN-PINN].
In this work, we introduce a memory-inspired, non-linear quantum neural network for entanglement engineering that combines physically motivated non-linear response functions with photonic beam-splitter–based architectures in a variational optimization framework. Upon training, both the optimized parameters and network topology can be directly mapped onto experimentally realizable photonic setups.
In Section II, we introduce the entanglement metrics, non-linear response functions, gate decomposition, and noise models used for NISQ simulations. Section III presents the numerical results, and Section V summarizes the conclusions and outlook.
II Methods
@@
Quantifying the degree and structure of entanglement in many-body quantum systems is essential for understanding their computational and physical capabilities [PRL-multipartiteEntanglementWitnesses, PRA-multipartyEntWintessForQuantumComputing, Review_multiPartyEnt, CV_multyPArtyEnt]. The core idea here is to realize entangled many-body states using a non-linear QNN, built using only one type of gate. The problem is thus defined as an entanglement maximization – or equivalently, cost minus-sign-function minimization, for both pure and mixed states. For NISQ-type simulations, we use native dephasing and amplitude damping circuits provided by $\mathrm{PennyLane}$ [PennyLane].
@@
By construction, the measure satisfies $0\leq Q(\psi)\leq 1$, with one corresponding to a GHZ state. It vanishes if and only if the state is fully separable, while values approaching unity correspond to maximally entangled states.
II.2 The photonic Beam Splitter Gate
Generally, quantum neural networks are blocks of unitary gates trained by shifting the rotation angles $\theta_{i}$ for each gate, in analogy to a classical neuron. But quantum networks usually lack non-linear functions that modify the neuron’s output. In this work, to introduce non-linear, physically inspired behaviour, we force our angles to follow a non-linear function $\theta=f_{\rm non-linear}(t)$, for a fictitious time parameter. The gate is implemented natively in PennyLane and acts on a pair of qubits (or optical modes encoded in qubits), labeled $\rm A$ and $\rm B$. Its action is parametrized by a reflectivity function $R(t)$, which encodes the physical time-dependent response of the underlying memristive element.
$\{T_{\mathrm{osc}},\,T_{\mathrm{int}},\,a,\,b\},$ describing, respectively, the intrinsic oscillation period, the integration window of the memristor, and two phenomenological constants entering the reflectivity model.
Given the reflectivity value $R(t)$, the gate compute an effective beam-splitter angle:
$\theta(t)=2\arcsin\,\Big(\sqrt{1-R(t)}\Big)$ (2)
which determines the strength of the conditional interaction between the two modes n physical components. To compare with the SIREN neural network, we consider a generic reflectivity:
$\theta(t)=\sin(t).$ (3)
For the linear network, the rotation angle of the unitary gate is reduced here to $\theta=1\cdot t$
Gate decomposition – The unitary decomposition in native operations of the beam-splitter gate is guaranteed by the Solovay-Kitaev theorem [SolovayKitaevTheorem]; in this way, we can run them using standard quantum circuit simulators:
@@ II.2 The photonic Beam Splitter Gate
    A single-qubit rotation
    $\mathrm{RY}\,\left(\frac{4\pi t}{T_{\mathrm{osc}}}\right)\quad\text{acting on wire A,}$
    which introduces a controlled phase accumulation proportional to the evolution time.
@@
A final $\mathrm{SWAP}$ gate ensures symmetry between the modes and matches the canonical ordering of a two-port beam splitter. This entire block define an operation that we call from now on $\mathbf{U}^{\rm BM}$ for a gate that implements (2) as non-linear function, and $\mathbf{U}^{sin}$ for the one that implement (3), and in case of linear network we use standard notation $\mathbf{U}^{\theta}$; in general we refer to it as $\mathbf{U}^{\varkappa}$ where $\varkappa=\{\rm RM,sin,\theta\}$.
II.3 The architecture and optimization task
Throughout this work, we make use of the $\mathbf{U}^{\varkappa}$ only to build our networks $\mathcal{U}_{\rm...}$. Each network possesses a specific topology, which describes how each $\mathbf{U}^{\rm BM}$ connects to couples of qubits. So, a topology can simply be described as a list $\left[(A_{1},B_{1})\dots(A_{n},B_{n})\right]$ to apply to our $n$ gates. In this work, we make use of 2 key topologies, random (RN), staircase (SC):
SC $=\mathbf{U}^{\varkappa}_{0,1}\mathbf{U}^{\varkappa}_{1,2}\cdot\cdot\mathbf{U}^{\varkappa}_{n-2,n-1}$ (4)
RN $=\mathbf{U}^{\varkappa}_{u_{0},u_{1}}\mathbf{U}^{\varkappa}_{u_{1},u_{2}}\cdot\cdot\mathbf{U}^{\varkappa}_{r_{n-1},r_{n}},$ (5)
@@ II.3 The architecture and optimization task
So, e.g., a network we call a network with SC topology that implements (2) as $\mathcal{U}^{\rm BM}_{SC}=\bigotimes\mathbf{U}^{\rm BM}_{0,1}\cdots\mathbf{U}^{\rm BM}_{n-2,n-1}$. We will also consider combinations of these networks, defined when needed.
With this in hand, we can now define our entanglement generation problem as a minimization one as follows:
@@
for any possible topology hereafter considered. Importantly, for the mixed state, we still employ the MW, because it is a scalable, experimentally accessible surrogate cost function, whose effectiveness we validate against negativity as a certification metric.
II.4 Noisy simulation
To depict a more realistic scenario, it is necessary to insert sources of noise at each gate level of our simulated photonic hardware. For this reason, we implement standard hardware noise sources, namely a dephasing channel after the $\mathrm{RY}$, and an amplitude-damping after each wire of the simulated memristor. The noisy channel is defined as
@@ II.4 Noisy simulation
with $Z$ the standard Pauli operator, and the Krauss matrices for the amplitude damping channel given by
@@
where $\|\cdot\|_{1}$ denotes the trace norm.
III Results
A complete analysis for the noiseless pure-states scenario is insightful, and so is offered in the Appendix B.1. We briefly state its key insights: (i) from the Monte Carlo analysis of $10^{5}$ random circuits, we show that the use of non-linear functions offers a significant boost for the entanglement production, and (ii) we can scale up to 20 qubits simply by implementing the SC topology, see Fig.12.
IV 5 qubits on low noise analysis.
@@ III Results
Figure 1: Meyer-Wallach and bipartite negativity for the $\mathcal{U}_{\rm sc}^{\rm BM/sin}$ models, i.e, the simple staircase model, respectively for the upper and lower panel. We can see that the variational optimization of this topology has poor optimization and produces a low amount of bipartite mixed state entanglement, for both bipartitions. This clearly prompts us towards introducing more gates between non-nearest neighbor qubits.
Figure 2: Run one of the Meyer-Wallach maximization and negativity evolution for the sub-systems for a 5-qubit mixed state, with only amplitude damping noise with $\gamma=0.01$. Upper panel, MW and negativity evolution for the $\mathcal{U}^{\rm BM}$. Lower panel, the same evolution but for the $\mathcal{U}^{\rm sin}$ topology.
@@
Figure 3: Another run for different random weights initialization of the networks $\mathcal{U}_{\rm sc+1}^{\rm BM/sin}$ . In the upper panel, we can clearly see how the global minima is quickly found for moderately noisy data with this random initialization, while this is not the case for the $\mathcal{U}_{\rm sc+1}^{\rm sin}$ model.
By comparing Fig.2 with Fig.3, we can notice the dependency on the initial condition. $\mathcal{U}^{\rm BM}_{\rm sc+1}$ is bound to achieve greater performance, a value of $\sim 1.4$ when the theoretical upper bound is $1.47$ bits, see the upper panel in Fig.3, but the dependency from network initialization conditions is likely. However, both architectures can get stuck in local minima, while this is not the case for the noiseless simulations.
IV.1 NISQ scenario
@@
As stated in Sec.II, the MW metric is now used as a valid cost function and scalable surrogate, and the multi-partite entanglement is certified with bipartite negativity. We begin by considering five-qubit networks with $U^{\mathrm{SC+1}}_{\mathrm{BM/sin}}$ topology, with a final coupling between qubits $(A,B)=(4,0)$. To assess the impact of circuit connectivity, we also introduce an enhanced topology denoted
$U^{\mathrm{SC+2}}_{\mathrm{BM/sin}}=U^{\mathrm{SC+1}}_{\mathrm{BM/sin}}\otimes U^{\mathrm{BM/sin}}_{3,1},$
@@
By contrast, the $\mathrm{SC+2}$ topology significantly improves both convergence and robustness. As illustrated in Fig.5, the additional coupling enables smoother optimization dynamics and leads to comparable negativity values across complementary bipartitions. In particular, for the sinusoidal non-linear activation, the achieved negativity approaches the corresponding theoretical upper bound for the considered partitions, indicating the effective generation of multipartite entanglement despite the presence of noise.
@@
Figure 5: Again, full noisy simulation with dephasing channel after $\mathrm{RY}$ and amplitude damping after $U_{\varkappa}$. Upper panel and lower panels showing the BM and sin simulations as usual. Two important observations follow: (i) the $\mathcal{U}_{\rm sc+2}^{\varkappa}$ topology has an extremely smoother training, with both sub-partitions achieving quickly the same value of negativity, and (ii) for the sinesine the theoretical upper bound is very close.
@@
In Fig.5 we can appreciate the effects of this topology, which stitches together the opposite qubits of our many-body state. Not only is the training smooth and faster, but, as can be noticed in the lower panel for $\mathcal{U}_{\rm sc+2}^{\rm sin}$, the two bipartitions can achieve a value of negativity which is close to the theoretical maximum of $\sim 1.47$.
IV.2 10 qubits noisy. Topology study
Last, we scale the problem for a many-body state of 10 qubits. Using the insights offered from the previous simulations, we test three types of networks for sinesine only non-linear term:
@@
$\mathcal{U}_{10}^{[0-3]}$ $=\mathcal{U}_{\rm sc+1}\mathbf{8,1}\mathbf{U}_{7,2}\mathbf{U}_{6,3}\mathbf{U}_{4,1}\mathbf{U}_{5,2}$
@@
In Fig.6, we can at what extent the topology design can influence the production of entanglement between bipartitions. In the plot (a), we can see the action of $\mathcal{U}_{10}^{[5-9]}$. Our engineering solution is well-suited for symmetric bi-partition. Forcing the entanglement into B= [0,1,2,3], instead, is more difficult. Comparing panel (b) with (c), we can see how the network $\mathcal{U}_{10}^{[0-3]}$ and $\mathcal{W}_{10}^{[0-3]}$ redistribute it. What is worth to highlight is that $\mathcal{U}_{10}^{[0-3]}$, in panel (b), has fewer gates, 12 instead of 14, but has better performance, obtaining $\sim 6$ of negativity against the $\sim 5.6$ from $\mathcal{W}_{10}^{[0-3]}$. So, for non-symmetric bi-partitions, increasing the number of gates does not seem to be a fruitful path.
@@
We have introduced a low-depth, linearly scaling quantum neural network architecture for variational entanglement engineering, inspired by memory-enabled photonic components and incorporating physically motivated non-linear activation functions. By combining non-linearity with simple two-qubit beam-splitter–based operations, and using the Meyer-Wallach as a surrogate cost function, the proposed framework enhances the expressive power of variational quantum circuits without increasing circuit depth beyond linear scaling.
@@
Through numerical simulations and Monte Carlo sampling over circuit topologies, we identified network structures capable of efficiently generating highly entangled pure states, approaching the GHZ limit, using several two-qubit operations proportional to the system size. These results highlight the central role of non-linearity and topology in enabling scalable multipartite entanglement generation.
@@
We further investigated the performance of the architecture in realistic noisy scenarios. In this regime, we certify multipartite entanglement using bipartite negativity. For mixed states of up to ten qubits, the optimized circuits produce substantial entanglement across both symmetric and asymmetric bipartitions. With the symmetric case achieving a negativity of 10 for a theoretical bound of $\sim 15$, while for the asymmetric case, the problem is more complex, and we can achieve a $\sim 6$ in entanglement production. Importantly, increasing the number of gates turns out as a less efficient route. This is a striking difference compared to the noiseless scenario, where increasing the number of $\mathrm{CNOT}$ and $\mathrm{SWAP}$ gates is bound to improve the amount of state entanglement. Taken together, our results establish a lightweight and experimentally motivated variational framework for engineering multipartite entanglement on near-term photonic platforms, in both pure and noisy quantum systems. Future endeavors will be devoted to a higher in-depth understanding, likely implementing also reinforcement learning or agentic solutions, together with a systematic effort for its theoretical description.
VI acknowledgment

### v2

# A nonlinear quantum neural network framework for entanglement engineering

@@ Abstract
Multipartite entanglement is a key resource for quantum technologies, yet its scalable generation in noisy quantum devices remains challenging. Here, we propose a low-depth quantum neural network architecture with linear scaling with a novel approach to introducing activation functions, used for the task of entanglement engineering. As a testbed to demonstrate the clear advantage unlocked by the introduction of nonlinear activations, we run a Monte Carlo sampling over $10^{5}$ circuit topologies for pure noiseless states. Supported by the clear edge offered by our approach, we focus our attention on the noisy scenario; we employ the experimentally accessible Meyer–Wallach global entanglement as a scalable surrogate optimization cost and certify entanglement using bipartite negativity. For mixed states of up to ten qubits, the optimized circuits generate substantial entanglement across both symmetric and asymmetric bipartitions. These results establish an experimentally motivated and scalable framework for engineering multipartite entanglement on near-term quantum devices, highlighting the combined role of nonlinearity and circuit topology that can achieve up to 20 qubits for very low noise levels.
@@ I Introduction
Entanglement engineering can be formally cast as a graph-optimization problem, in which both interaction structure and parametrization determine the entanglement properties of the resulting quantum state [QuantumExperimentsAndGraphs, EntanglementGenerationKobra]. This perspective has motivated the use of machine-learning techniques, particularly optimization and reinforcement learning, in quantum information science [Review_AIforQauntum-Krenn, ReviewAIforQuantumComp-2025, RL_subgraphOptimization]. Such approaches have been successfully applied to the automated design of quantum experiments [VAE_EntanglementExperiments-2022], circuit synthesis [RL4qEntangledQuibits, DiffusionCircuitsynthesis-2024], circuit optimization [RL_circuitOPtimization], and the preparation and engineering of entangled states [RLStatePReparation, RLEntanglementStatePreparation, RL_entanglementEgineering].
@@
Quantum neural networks (QNNs) provide a flexible framework for these tasks by optimizing parametrized quantum circuits using classical learning algorithms [schuld2021supervisedquantummachinelearning]. However, most existing QNN architectures rely on strictly linear parametrizations, which limits their expressive power at fixed circuit depth. Increasing depth can partially alleviate this limitation but leads to higher hardware noise and reduced scalability, making low-depth, linearly scaling architectures particularly attractive for near-term devices [Luo_2023].
A key ingredient of our approach is the introduction of different nonlinear behaviors. While nonlinear activation functions are essential for the success of classical neural networks, only a few quantum architectures incorporating nonlinearity have been explored so far [QCNN-nonlinear, Reservoir-Spagnolo2022, CVNonlinearQNN, AlbertoMemristor, TripartiteEntMemristor, EntangledQuantumMemristor]. This represents a missing yet important step in statistical learning, whose role is well established [Activactoins-Kunc2024]. The practical problem is that, while implementing an activation function is straightforward for classical networks, it would demand the realization of specific hardware gates and experimental solutions for a quantum network.
In this work, we address the challenge of engineering multipartite entanglement at fixed circuit depth under realistic noise by introducing a quantum neural network architecture with memory-inspired nonlinear parameter response. Rather than increasing depth or gate count, our approach exploits nonlinearity and circuit topology as complementary design resources to enhance the probability of generating highly entangled states.
To demonstrate the edge offered by our novel application of nonlinear functions, we run a Monte Carlo sampling over circuit architectures to showcase it right off from the noiseless case. To enable scalable optimization for mixed states, we employ the Meyer–Wallach global entanglement as a computationally efficient surrogate cost function, and certify entanglement using bipartite negativity. Our results establish a principled and experimentally motivated framework for multipartite entanglement engineering on near-term quantum devices, highlighting nonlinearity as a key ingredient for noisy algorithms.
We first demonstrate the edge offered by the presence of a nonlinear term, focusing mostly on simulations where the presence of noise is relevant. Our solution can be used in two ways: (i) use the training weights after the training, or (ii) directly realize it on photonic setups.
The paper is organized as follows: in Section II, we introduce the entanglement metrics, the nonlinear functions and their justification, and the gate decomposition for circuit situations. In Section III we define the model architecture, the training, and the metrics and training for the noisy scenario. Lastly, the noise gate simulates a Noise-Intermediate-Scale-Quantum (NISQ) scenario. Section IV presents the numerical results. Firstly, we offer an analysis for a pure state that demonstrates the clear boost offered by nonlinear functions. Then, we focus solely on the noisy scenario. Lastly, in Section V summarizes the conclusions and outlook.
II Preliminaries
@@
Quantifying the degree and structure of entanglement in many-body quantum systems is essential for understanding their computational and physical capabilities [PRL-multipartiteEntanglementWitnesses, PRA-multipartyEntWintessForQuantumComputing, Review_multiPartyEnt, CV_multyPArtyEnt]. The core idea here is to realize entangled many-body states using a nonlinear QNN, built using only one type of gate. We begin by defining the network’s building blocks: the key metric for the cost function, the new activation function strategy, and the only gate we use, which is a single photon beam-splitter-inspired gate for all our architectures. Lastly, we introduce the metric and further considerations for the NISQ scenario.
@@
By construction, the measure satisfies $0\leq Q(\psi)\leq 1$, with one corresponding to a maximally entangled state like the GHZ, but not a W-type state, for example. As expected, the metric vanishes if and only if the state is fully separable.
II.2 Nonlinear Control Reparameterization
Generally, quantum neural networks are blocks of unitary gates trained by shifting the rotation angles $\theta_{i}$ for each gate. Like their classical counterpart, they combine several elementary operations acting upon different qubits to realize different architectures. What is missing with respect to the classical counterpart is the implementation of nonlinear activation functions. This is a concrete drawback and lies behind the great success of deep learning; therefore, its lack is a drawback in quantum machine learning. For example, a multilayer perception (MLP) reduces to a combination of matrix multiplications plus nonlinear functions right after each of them, more precisely:
$\text{\rm MLP}(\bar{\theta})[\mathbf{x}_{\rm in}]=(\sigma_{2}\circ W_{2}^{\theta_{2}}\circ\sigma_{1}\circ W_{1}^{\theta_{1}})\mathbf{x}_{\rm in},$ (2)
with $\sigma_{i}$ a generic nonlinear function, $\mathbf{x}_{\rm in}$ a generic input, for a network optimizing on the set of weights parameter $\bar{\theta}$. In practical terms, they remap the neuron’s linear output into another space. While this can be implemented relatively simply in software, realizing a physical quantum gate that implements it is an open challenge. To tackle this setback, here we study another way around, oppositely implementing them:
[…]
$\tilde{\theta}$ $=\sigma(\theta).$
As Eq.  (3) states, the activation function now modifies the output forcing it inside its dominium, and is later on linearly used by a standard gate. The first $\sigma$ we test is inspired by a photonic quantum memristor with single photon input [AlbertoMemristor]. Given a physical-inspired reflectivity value $R(\theta)$, see Eq.  (14), the gate simulates an effective beam-splitter angle:
$\tilde{\theta}=2\arcsin\,\Big(\sqrt{1-R(\theta)}\Big),$ (4)
which, in physical settings, determines the strength of the conditional interaction between the two modes of physical components. What is physically interesting is that this type of function brings the hallmark of quantum memory effects unraveled by a memristor-type beam-splitter. In other words, even though the gate is not implemented as a genuine memristor component, it is optimized as such.
Following Eq.  (14) and Eq.  (4), we choose as the second $\sigma_{\rm sin}$. This type of function draws inspiration from a recent breakthrough in the classical field of deep learning, dubbed Sinusoidal Representation Network (SIREN)[mildenhall2022neuralfields, SIREN, SIREN-PINN, HSIREN]. Therefore, we also implement the SIREN-type nonlinearity:
$\tilde{\theta}=\sin(\theta).$ (5)
Last, to benchmark the advantage offered by the method, we consider the unitary function $\bar{\theta}=1\cdot\theta$.
Photonic quantum memristor gate decomposition – The unitary decomposition in native operations of the photonic quantum memristor (PQM) offered in [AlbertoMemristor], when restricted to only two can be decomposed using standard quantum circuit [SolovayKitaevTheorem]:
@@ II.2 Nonlinear Control Reparameterization
    A single-qubit rotation $\mathrm{RY}$ acting on wire A, which introduces a controlled phase accumulation proportional to the evolution time.
@@
  3. 3.
    Last, a final $\mathrm{SWAP}$ gate ensures symmetry between the modes and matches the canonical ordering of a two-port beam splitter.
This entire block simulates the behaviour of a memristor-type beam splitter, as the one studied in Ref. [AlbertoMemristor]. From now on, we call this block of operations $\mathbf{U}^{\rm BM}$ for a gate that implements Eq. (4) as nonlinear function, and $\mathbf{U}^{sin}$ for the one that implement Eq. (5), and in case of linear network we use standard notation $\mathbf{U}^{\theta}$; in general we refer to it as $\mathbf{U}^{\varkappa}$ where $\varkappa=\{\rm RM,sin,\theta\}$.
III Methods
III.1 The architecture and optimization task
Throughout this work, we make use of the $\mathbf{U}^{\varkappa}$ only to build our networks, which, in general, we define with the symbol $\mathcal{U}_{\rm topology}^{\varkappa}$, whose definition we introduce below. Each network possesses a specific topology, which describes how each $\mathbf{U}^{\rm BM}$ operation (or gate) connects to couples of qubits. So, a network topology can simply be described as a list $\left[(A_{1},B_{1})\dots(A_{n},B_{n})\right]$ to apply to our $n$ gates. In this work, we make use of 2 key topologies, random (RN), staircase (SC):
SC $=\mathbf{U}^{\varkappa}_{0,1}\mathbf{U}^{\varkappa}_{1,2}\cdot\cdot\cdot\mathbf{U}^{\varkappa}_{n-2,n-1}$ (6)
RN $=\mathbf{U}^{\varkappa}_{u_{0},u_{1}}\mathbf{U}^{\varkappa}_{u_{1},u_{2}}\cdot\cdot\cdot\mathbf{U}^{\varkappa}_{r_{n-1},r_{n}},$ (7)
@@ III.1 The architecture and optimization task
So, e.g., a network with SC topology that implements the beam-splitter type nonlinearity Eq. (4) will be written as $\mathcal{U}^{\rm BM}_{SC}=\bigotimes\mathbf{U}^{\rm BM}_{0,1}\cdots\mathbf{U}^{\rm BM}_{n-2,n-1}$. We will also consider combinations of these networks, defined when needed.
With this in hand, we can now define our entanglement generation problem as a minimization problem as follows:
@@
for any possible topology hereafter considered. Importantly, for the mixed state, we still employ the MW because it is a scalable, experimentally accessible surrogate cost function, whose effectiveness we validate against negativity as a certification metric. Last, it is worth stressing the analogy with the classical counterpart; similarly to SIREN networks, a single input comprises the entire dataset.
III.2 Noisy simulation. Implementation and validation metric
To depict a more realistic scenario, it is necessary to insert noise sources at each gate level of our simulated photonic hardware. For this reason, we implement standard hardware noise sources, namely a dephasing channel after the $\mathrm{RY}$, and an amplitude-damping after each wire of the simulated memristor. The noisy channel is defined as
@@
with $Z$ the standard Pauli operator, and the Kraus matrices for the amplitude damping channel given by
@@
where $\|\cdot\|_{1}$ denotes the trace norm. Relative to the MW global entanglement, this is no longer interpreted as a strict entanglement monotone and an entanglement validation metric, but rather as a scalable proxy and a good surrogate cost function, promoting multipartite correlations during network optimization.
Training hyperparameters – For the optimization, we implement the classical Adam optimization, a learning rate of $5\cdot 10^{-3}$, and early stopping for regularization. The max circuit depth considered is two. $\mathrm{Pennylane}$ is used to code and train all the models  [PennyLane].
IV Results
The first mandatory question to assess is the boost offered by nonlinearities inside our optimization scheme. A Monte Carlo analysis of this key question for the noiseless pure-states scenario is offered. The main outcomes hereafter displayed will leverage the numerical evidence and focus on a step-by-step analysis, starting from 5 low-noise qubits up to 10 fully noisy qubits.
Activation functions advantage – The histograms in Fig.  6 in the Appendix .1 display the validity of our hypothesis: (i) from the Monte Carlo analysis of $10^{5}$ random noiseless circuits, we show that the use of nonlinear functions offers a significant boost for the entanglement production, and (ii) in this regime, because the Meyer-Wallach is both an entanglement cost function and a valid validation metric, we can scale up to 20 qubits simply by implementing the SC topology for almost pure states, see Fig. 11.
IV.1 5 qubits on low noise analysis
@@ IV Results
Figure 1: Upper panel. Meyer-Wallach and bipartite negativity for a single run of the $\mathcal{U}_{\rm sc}^{\rm BM}$ models, i.e, the simple staircase model. Lower panel, the same plot for $\mathcal{U}_{\rm sc}^{\rm sin}$ model. We can see that the optimization of this topology produces an amount of bipartite mixed state entanglement $\sim 0.5$, for both bipartitions, also in the presence of very low noise. This clearly prompts us to introduce more gates between nonnearest-neighbor qubits.
@@
Figure 2: Upper panel. Mean value and standard deviation training evolution for 20 different random initializations of the network $\mathcal{U}_{\rm sc+1}^{BM}$. Lower panel. The same for $\mathcal{U}_{\rm sc+1}^{sin}$ networks. In the upper panel, we can see that the memristor-inspired architecture achieves a higher standard deviation, but can potentially obtain a close-to-theoretical value for B=[0,1,2], while this is not the case for the $\mathcal{U}_{\rm sc+1}^{\rm sin}$ model. This shows the potential and the sensitivity as well to the network’s initialization values.
Undoubtedly, Fig. 2 confirms the sensitivity of the network to initial parameter weights. Looking at the upper panel, we can see that $\mathcal{U}^{\rm BM}_{\rm sc+1}$ is bound to achieve greater performance, a value of $\sim 1.46$ when the theoretical upper bound is $1.5$, demonstrating the efficiency of the approach for a small number of qubits. We can appreciate now how the $\mathcal{U}_{\rm sc+1}^{\rm BM/sin}$ models’ standard deviation in the learning curve is evident now, of the order of $\sim 10^{-1}$.
IV.2 5 qubits NISQ scenario
@@
As stated in Sec.  III, the MW metric is now used as a valid cost function and scalable surrogate, and the multi-partite entanglement is certified with bipartite negativity. We begin by considering five-qubit networks with $U^{\mathrm{SC+1}}_{\mathrm{BM/sin}}$ topology, with a final coupling between qubits $(A,B)=(4,0)$. To assess the impact of circuit connectivity, we also introduce an enhanced topology denoted
$\mathcal{U}_{\mathrm{SC+2}}^{\mathrm{BM/sin}}=\mathcal{U}_{\mathrm{SC+1}}^{\mathrm{BM/sin}}\otimes\mathbf{U}^{\mathrm{BM/sin}}_{3,1},$ (13)
@@
By contrast, the $\mathrm{SC+2}$ topology significantly improves both convergence and robustness. As illustrated in Fig.4, the additional coupling enables smoother optimization dynamics and leads to comparable negativity values across complementary bipartitions. In particular, for the sinusoidal nonlinear activation, the achieved negativity approaches the corresponding theoretical upper bound for the considered partitions, indicating the effective generation of multipartite entanglement despite the presence of noise.
@@
Figure 4: Mean value and standard deviation of the evolution of the full noisy simulation with dephasing channel after $\mathrm{RY}$ and amplitude damping after $U_{\varkappa}$, for 20 different random initialization of the network. Upper panel and lower panels showing the BM and sin simulations as usual. We can observe that the $\mathcal{U}_{\rm sc+2}^{\varkappa}$ can potentially achieve close to the theoretical value, and the physical memristor-inspired $\mathcal{U}_{\rm sc+2}^{\rm BM}$ is bound to be slightly more efficient. This confirms the sensitivity to the network’s parameter initialization values, but different from Fig.  2, this now reduces, as the difference in standard deviation demonstrates. This can be interpreted as a positive side effect of noise injection inside the differential landscape.
@@
In Fig. 4 we can appreciate the effects of this topology in Eq.  (13), whose structure is meant to stitch together the opposite qubits of our many-body state. Not only is the training smooth and faster, but, as can be noticed in the lower panel for $\mathcal{U}_{\rm sc+2}^{\rm sin}$, the two bipartitions can achieve a value of negativity which is close to the theoretical maximum of $\sim 1.5$.
IV.3 10 qubits NISQ scenario. Topology study
Last, we scale the problem for a many-body state of 10 qubits. We opt for testing three types of networks for sinesine only nonlinear term:
@@
$\mathcal{U}_{10}^{[0-3]}$ $=\mathcal{U}_{\rm sc+1}\mathbf{U}_{8,1}\mathbf{U}_{7,2}\mathbf{U}_{6,3}\mathbf{U}_{4,1}\mathbf{U}_{5,2}$
@@
In Fig. 5, we can at what extent the topology design can influence the production of entanglement between bipartitions. In the plot (a), we can see the action of $\mathcal{U}_{10}^{[5-9]}$. Our engineering approach is particularly effective for symmetric bi-partition. Forcing the entanglement into B= [0,1,2,3], instead, is more difficult. Comparing panel (b) with (c), we can see how the network $\mathcal{U}_{10}^{[0-3]}$ and $\mathcal{W}_{10}^{[0-3]}$ redistribute it. It is worth to highlight is that $\mathcal{U}_{10}^{[0-3]}$, in panel (b), has fewer gates, 12 instead of 14, but has better performance, obtaining $\sim 6$ of negativity against the $\sim 5.6$ from $\mathcal{W}_{10}^{[0-3]}$. So, for non-symmetric bi-partitions, increasing the number of gates does not seem to be a fruitful path.
@@
We have introduced a low-depth, linearly scaling quantum neural network architecture for entanglement engineering, inspired by memory-enabled photonic components and incorporating physically motivated nonlinear activation functions. By combining nonlinearity with simple two-qubit beam-splitter–based operations and using the Meyer-Wallach as a surrogate cost function, the proposed framework enhances the expressive power of quantum circuits without increasing circuit depth beyond linear scaling.
@@
Through numerical simulations and Monte Carlo sampling over circuit topologies, we identified network structures capable of efficiently generating highly entangled pure states, using several two-qubit operations proportional to the system size. These results highlight the central role of nonlinearity and topology in enabling scalable multipartite entanglement generation.
@@
We further investigated the performance of the architecture in realistic noisy scenarios. In this regime, we certify multipartite entanglement using bipartite negativity. For mixed states of up to ten qubits, the optimized circuits produce substantial entanglement across both symmetric and asymmetric bipartitions. With the symmetric case achieving a negativity of 10 for a theoretical bound of $\sim 15$, while for the asymmetric case, the problem is more complex, and we can achieve a $\sim 6$ in entanglement production. Importantly, increasing the number of gates turns out as a less efficient route. This is a striking difference compared to the noiseless scenario, where increasing the number of $\mathrm{CNOT}$ and $\mathrm{SWAP}$ gates is bound to improve the amount of state entanglement. Taken together, our results establish a lightweight and experimentally motivated framework for engineering multipartite entanglement on near-term photonic platforms, in both pure and noisy quantum systems. Future endeavors will be devoted to a higher in-depth understanding, likely implementing also reinforcement learning or agentic solutions, together with a systematic effort for its theoretical description.
Acknowledgements.